Oxygen gas expands reversibly in a cylinder behind a piston at a constant pressure of 3 bar. The initial and final volumes are 0.01m3 and 0.03m3 respectively.The initial temperature is 17 degrees celsius. Calculate ,(a)the work input. (b)the heat supplied during the expansion. Assume oxygen to be a perfect gas with molar mass 32kg/mol and take Cp=0.917 kj/kgK


1
Expert's answer
2021-05-19T06:07:38-0400

a)


"W=p(V_f-V_i)\\W=3\cdot10^5(0.03-0.01)=6000\ J"

b)


"3\cdot10^5(0.01)=\frac{m}{32}(8.31)(290)\\m=39.8165\ g\\n=\frac{39.8165}{32}=1.2443\ mol\\3\cdot10^5(0.03)=\frac{m}{32}(8.31)T_2\\T_2=870\ K\\Q=(39.81)(0.917)(870-290)=\ 21180\ J"